If a square has a diagonal that measures 12, and so the square is divided into 2 triangles, what is the value of one of the sides of the square?
Answer by mukhopadhyay(490)

Let each side of the square is x unit;
x^2+x^2 = (12)^2 = 144
=> x^2 = 72
=> x = 6 sqrt(2)
Length of each side of the square is 6 Sqrt(2) unit